What quantity of 60 per cent acid solution must be mixed with a 25 percent solution to produce 252 mL of a 50 per cent solution?
Answer by Alan3354(69443)

What quantity of 60 per cent acid solution must be mixed with a 25 percent solution to produce 252 mL of a 50 per cent solution?
-------------
s = 60%
t = 25%
s+t = 252 (amount of solution)
0.6s + 0.25t = 0.5*252 (amount of acid)
t = 252-s
Sub for t
0.6s + 0.25(252-s) = 126
0.6s + 63 - 0.25s = 126
0.35s = 63
s = 180 ml
t = 72 ml